OpenStudy (anonymous):

Write the standard equation for the circle that passes through the points: (-1, 2) (4, 2) (- 3, 4) has to be in x^2 + y^2 + dx + ey + f = 0 form plz help Idk how to do this at all

OpenStudy (anonymous):

1+4-d+4e+f=0 16+4+4d+2e+f=0 9+16-3d+4e+f=0

OpenStudy (anonymous):

^^^simplified(-d+4e+f=-5) (4d+2e+f=-20) (-3d+4e+f=-25)

OpenStudy (dumbcow):

solve system of 3 equations first eliminate one of the variables thus making it a system of 2 equations then eliminate another variable, leaving a single equation solve for remaining variable then use back substitution to obtain other variables

OpenStudy (dumbcow):

start with "f" first 4d +2e +f = -20 -(-d +4e+f = -5 ) ------------------------- 5d -2e = -15 4d +2e +f = -20 -(-3d+4e+f=-25) ------------------------- 7d -2e = 5

OpenStudy (dumbcow):

now eliminate "e" 5d -2e = -15 -(7d -2e =5) ----------------- -2d = -20 ---> d = 10

OpenStudy (anonymous):

okay so how would i find e and f?

OpenStudy (dumbcow):

substitution

OpenStudy (dumbcow):